OpenStudy (anonymous):

what is the derivative of y= (x+1)^(13)?

OpenStudy (anonymous):

do you know the power rule y=x^n y'=n*x^(n-1) also the chain rule is used here: y = F(g(x)) so y'= f '(g(x))*(g'(x))

OpenStudy (anonymous):

dy/dx = 13(x+1)^12 * d/dx(x+2) = 13(x+1)^12

OpenStudy (anonymous):

...yes if you just want to give the answer that is it.

OpenStudy (anonymous):

and if i'll solve for the 13th derivative, i'll just do that on and on right?

OpenStudy (anonymous):

yes since the inside function is x with a constant, the derivative every time will be d/dx (x+1) which is one so the second derivative will be 13(x+1)^12 = 13*12(x+1)^11

OpenStudy (anonymous):

what is the derivative of y= sqrt of (2x)?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

too complex to type lol

OpenStudy (anonymous):

i will explain it in words..

OpenStudy (anonymous):

isn't it